May is tornado month in Oklahoma. According to the National Weather Service, on average in May, the number of tornadoes in Oklahoma is 21.7 tornados. Consider the random variable X the amount of time (in days) until the 10th tornado occurs.HINT: There are 31 days in May. Think about, on average, a tornado occurs every ___ days.

What is the standard deviation of the number of days before the 10th tornado?
Answer #1

on average, a tornado occurs every 21.7/31=0.7 days

standard deviation of the number of days before the 10th tornado =sqrt(10*0.72) =2.2136 Days (as it follows gamma distribution with parameter =10 and =0.7)